# Machine learning surrogates for efficient hydrologic modeling: Insights from stochastic simulations of managed aquifer recharge

### v1

@@ 2.5 Data preprocessing
To train each ML surrogate model, we designate the inputs of the ParFlow-CLM simulations (e.g., 3D porosity fields, transient meteorological forcing, etc.) as our training features and the outputs of the ParFlow-CLM simulations (transient, 3D pressure fields) as our training targets. However, prior to training, we first perform several steps of preprocessing. In particular, we perform feature selection to curate a list of salient input features, we broadcast input features across the $x$, $y$, $z$ and $time$ dimensions to standardize input and output formats, and we downsample certain inputs to reduce memory requirements for training.
@@
Field name Units $\mu$ $\sigma$ simulations variable variable 1 2 3
@@
    At the start of model Stage 1 (model spin up to steady state), pressure across the domain is in hydrostatic equilibrium with the water table at 45 m depth. For Stages 2 and 3, the initial pressure head field is initialized as the final pressure field output from the previous stage.
@@ 2.5.1 Feature selection
We select 17 input fields as the input features and one output field as our supervised learning target (Table 2.5). These input fields include hydrogeologic parameters (fields 1-6 in Table 2.5), meteorologic forcing (fields 7-14), information about the rectilinear model grid (field 15), initial conditions (field 16) and the recharge forcing applied at the surface (field 17).
The process-based hydrologic model receives additional input information not included in Table 2.5, such as cell dimensions and time step sizes. However, we omit many of these fields from our input features because they are temporally invariant and constant across all training examples. Thus, each surrogate architecture likely learns this information implicitly during training. Some fields, such as cell volume and the initial pressure head for Stage 1, are constant across all simulations, but are included nonetheless to support future studies seeking to modify these parameters.
@@ 2.5.2 Broadcasting
Broadcasting streamlines manipulation of inputs and targets through the use of large, combined tensors. The input features (Table 2.5) include fields that remain constant in time, fields that constant in space and fields that are constant in both space and time. Broadcasting simply repeats the values of each field across dimensions where that field remains constant, which ensures that the tensor for each input feature is four-dimensional (x, y, z and time). For example, we broadcast air temperature at each time step — which does not vary spatially across the domain — along the x, y and z dimensions. Similarly, given that saturated hydraulic conductivity does not change from one time step to the next, we repeat the 3D hydraulic conductivity field from the first time step across all time steps. This approach standardizes the dimensions of all input fields and allows them to be combined into one large tensor for input into each architecture.
@@ 2.5.3 Downsampling
While broadcasting standardizes input dimensions, it increases the total size of the input features and, by extension, the memory allocated to each training example. The memory required to train an ML model is an important component of computational efficiency; if the input and output for a single training example exceeds the memory capacity of a single graphics processing unit (GPU), then the model will have to be trained across multiple GPUs, which will increase the total number of processor-hours used during training. Modern hydrologic model domains can include anywhere from $10^{5}$ to $>$$10^{8}$ total grid cells (Wood et al., 2011; Bierkens et al., 2015). At this scale, after broadcasting each input feature across the x, y, z and time dimensions and adding in the size of the training targets (i.e., the transient 3D pressure fields from ParFlow-CLM), the memory allocated to a single training example can quickly exceed 100 GB, which is larger than the memory capacity of most modern GPUs. To address this challenge, we limit the size of a single training example through temporal and spatial downsampling.
For model stages 1 and 2, we apply temporal downsampling. These model stages are only used to initialize water content throughout the vadose zone. Because intermediate pressure fields are not used to calculate recharge efficiency, decreasing the temporal resolution does not limit downstream analyses. For each stage, we sample 16 output time steps from the spin up period using a geometric sampling scheme:
$\left\{\left\lfloor\sum_{k=1}^{t}r^{k-1}\right\rfloor:\sum_{k=1}^{16}r^{k-1}=160\right\}_{t=1}^{16}.$ (1)
The length of 16 is chosen so that a single training example fits on our training hardware, while the geometric sampling scheme prioritizes time step outputs toward the beginning of the model stage, when changes in pressure head are most pronounced. Though the final steady-state pressure field at the end of each stage is the only output used in subsequent modeling stages, we calculate loss across multiple time steps from this stage to ensure that the model simulates physically meaningful processes (i.e., gradual changes in pressure over time). This also forces the model to output intermediate pressure fields; visually examining these intermediate outputs also allows the user to more easily diagnose factors contributing to poor model performance. While this geometric sampling scheme results in a variable time step size within an individual simulation, the output frequency is consistent across all training examples. Thus, the data-driven ML models should implicitly learn these temporal patterns during training.
@@
To downsample the Stage 3 training targets, we use an autoencoder. The autoencoder is a separate neural network that is trained to reproduce the spatially variable pressure field from a single time step, but it is forced to store this information in an intermediate, lower-dimensional space (Figure 3(a)). Each surrogate model is then trained on the lower-dimensional representation of the 3D pressure fields, known as the encoded targets (Figure 3(b)). We train a single autoencoder that is used for all surrogate models. To calculate surrogate performance on the test set, we then decode the surrogate model’s predictions using the autoencoder and calculate error metrics relative to the high-resolution pressure fields output by ParFlow-CLM. This workflow allows us to train our surrogate model in the reduced dimensional space, minimizing memory requirements, while retaining the high spatial resolution of our predictions for use in downstream tasks, such as calculating recharge efficiency. For a detailed overview of the autoencoder architecture, please refer to Section S1.3 in the Supplementary material.
@@ 2.6.1 Training configuration
where $y_{i}$ is the ground-truth output for the $i$th training example, $x_{i}$ is the input for the $i$th training example, $h(x_{i})$ is the predicted output, $w$ defines a depth-wise weighting scheme and $||\cdot||_{2}$ is the L2 norm. We choose this normalized loss function because it is particularly effective when there is large variance in the data and because previous hydrologic surrogate modeling work suggests that it provides faster convergence than the L1 loss (Wen et al., 2022). Given that managed aquifer recharge efficiency is largely governed by unsaturated flow processes (Perzan et al., 2023), accurately simulating pressure changes within the vadose zone is more important than simulating pressure changes deep within the saturated zone. Thus, we define the weights $w$ such that they more severely penalize pressure errors closer to the surface than at depth (Figure S3).
To investigate the impact of our choice of loss function on model performance, we train additional CNN3d surrogates on Stage 1 simulations using the mean square error loss (i.e., the square of the numerator of Equation 2). Due to computational constraints, we could not train duplicates of each surrogate architecture using the mean square error loss, so we only perform this comparison with the CNN3d because it has the fastest training time.
We perform a random 80%-10%-10% (217-28-28) training–validation–test split and all performance metrics are reported based on the test set unless otherwise noted. Features and labels are normalized using their respective minimum and maximum values from the training set. We normalize all splits by the training set’s minimums and maximums to prevent bias in validation and test results introduced by distribution information from one set to another. To assess the effects of data normalization on ML model accuracy, we train two additional Stage 1 CNN3d models: one using Z-score normalization and one without any data normalization. As discussed above, we only perform these tests using the CNN3d because of its short training time.
@@ 2.6.2 Hardware environment
We train all ML models on a single NVIDIA A100 Tensor Core graphics processing unit (GPU) with 80 GB memory. All ParFlow-CLM simulations were performed on AMD Epyc 7543 central processing units (CPUs). Depending on the model stage, ParFlow-CLM simulations were run either on 4 MPI ranks with 1 CPU core per rank (Stage 1 and Stage 3) or 6 MPI ranks with 1 CPU core per rank (Stage 2).
@@ 2.6.3 Model accuracy evaluation
We evaluate each surrogate model using the mean absolute percentage error (MAPE) between the predicted pressure head time-series and the ground-truth (ParFlow-CLM) pressure head time-series in a cell-by-cell manner. Because we train a different surrogate for each modeling stage, we report different MAPE values for each stage. As with the loss function, we weight MAPE by depth to ensure that each model’s performance accurately reflects its ability to simulate vadose zone flow (Figure S3). We also evaluate model performance when each surrogate is run sequentially across all three stages, such that the final pressure head output by the surrogate from Stage 1 is the initial pressure head (field 16 in Table 2.5) used for the Stage 2 surrogate model and the final pressure head output of Stage 2 is the initial pressure head for Stage 3. We call this technique end-to-end (E2E) evaluation. E2E provides a more realistic evaluation of surrogate performance in a real-world application.
Training an accurate ML surrogate requires a sufficiently large training set, but at a certain point, the cost of increasing the training set size may outweigh any benefit in improved prediction performance. To evaluate this relationship and estimate the minimum number of process-based simulations required to train an accurate surrogate, we train multiple versions of each Stage 1 surrogate model while using smaller and smaller training sets. In addition to the baseline training set size of 245 examples used for training (with a 217-28 training-validation split), we use training set sizes of 196 (174-22 split), 147 (130-17), 98 (87-11), 49 (43-6), 25 (22-3), 13 (11-2), 8 (7-1), 5 (4-1), and 3 (2-1) training examples while maintaining the test set constant at 28 examples. Because ML models typically exhibit a power-law relationship between model error and training set size (Hestness et al., 2017), we then fit a power function to these data ($y=am^{b}$, where $m$ is the number of training examples, $y$ is the test set MAPE and $a$ and $b$ are regression coefficients). Lastly, we differentiate $am^{b}$ with respect to $m$, which yields $abm^{b-1}$. This expression represents the instantaneous change in MAPE given an increase in training set size, which we use to compare each model’s sensitivity to training set size.
@@ 2.6.4 Model runtime evaluation
In addition to evaluating ML model accuracy, we quantify the total compute time needed to generate simulations with this hybrid modeling workflow. Since model runtime is largely dependent on the hardware on which it is executed, we strictly adhere to the hardware environment described in §2.6.2 when evaluating model runtime. To simplify analysis and minimize additional training, we only consider models trained on Stage 1 simulations in this section.
@@ 2.7 Summary of experiments
    We train all seven surrogate architectures — the CNN4d, ViT4d, U-FNO4d, CNN3d, ViT3d, U-FNO3d, and PredRNN++ (§2.4) — on Stage 1, 2, and 3 ParFlow-CLM simulations, amounting to 21 trained models. We also train an autoencoder for all Stage 3 models. This totals 22 trained models.
@@ 3.1 Model accuracy
All seven ML models achieve $<$10% MAPE on the test set (Figure 5 and Table S8), which is the threshold for “highly accurate forecasting” (Lewis, 1982). The PredRNN++ displays the lowest error for Stage 1 and end-to-end (E2E) predictions, while the CNN3d and ViT3d perform the best on Stage 2 and Stage 3, respectively. Even though the non-sequential models have more parameters than their one-step counterparts (Table 1), they do not exhibit an appreciable increase in predictive performance. The lone recurrent model (the PredRNN++), however, outperforms the others and is better able to learn the temporal dependencies that develop as each hydrologic simulation progresses through time. The large memory cells retained between time steps likely provide ample degrees of freedom for the model to store spatially heterogeneous information over time.
Figure 5: Test set MAPEs for all ML models across all stages (a) and for Stage 2, Stage 3 and end-to-end (E2E) predictions (b). The ML surrogate models are grouped by model type (see §2.4), with the non-sequential models on the left (CNN4d, ViT4d and U-FNO4d), the one-step models in the middle (CNN3d, ViT3d and U-FNO3d) and the recurrent model on the right (PredRNN++). Exact MAPE values are listed in Table S8.
Overall, prediction error is the highest for Stage 1, with a mean MAPE of 5.6% across all ML models, as opposed to 0.13% and 0.42% for Stages 2 and 3, respectively. Previous modeling work at this site has shown that managed aquifer recharge efficiency is most strongly controlled by vadose zone water content prior to inundation (Perzan et al., 2023). The low model error for Stages 2 and 3 makes sense, then, given that Stage 2 and 3 are provided the ground-truth pressure field output by ParFlow-CLM at the end of Stage 1 and 2, respectively, as initial pressure head fields during training.
Alternatively, the higher Stage 1 MAPE could result from its large time scale (§2.2 and 2.5.3). While Stage 2 and 3 span 1 and 2 years, respectively, Stage 1 spans over 400 years, with more dramatic changes in pressure between outputs. Pressure head changes across large time scales may be more difficult to learn than smaller, more incremental changes, potentially contributing to the poorer performance of Stage 1 models.

### v2

@@ 2.5 Data preprocessing
To train each ML surrogate model, we designate the inputs of the ParFlow-CLM simulations (e.g., 3D porosity fields, transient meteorological forcing, etc.) as our training features and the outputs of the ParFlow-CLM simulations (transient, 3D pressure fields) as our training targets. However, prior to training, we first perform several steps of preprocessing. In particular, we perform feature selection to curate a list of salient input features, we broadcast input features across the $x$, $y$, $z$, and time dimensions to standardize input and output formats, and we downsample certain inputs to reduce memory requirements for training.
@@
Field name Units $\mu$1 $\sigma$1 simulations variable variable 1 2 3
@@
    We provide the mean and standard deviation over all 273 simulations. Although not all fields follow a normal distribution, these metrics offer insights into the data’s range and variability.
  • 2
    At the start of model Stage 1 (model spin up to steady state), pressure across the domain is in hydrostatic equilibrium with the water table at 45 m depth. For Stages 2 and 3, the initial pressure head field is initialized as the final pressure field output from the previous stage. This reflects the fact that, when these simulations were generated, Stage 2 began simulations with the final state of Stage 1, and, similarly, Stage 3 began simulations with the final state of Stage 2.
@@ 2.5.1 Feature selection
We select 17 input fields as the input features and one output field as our supervised learning target (Table 2.5). These input fields include hydrogeologic parameters (fields 1-6 in Table 2.5), meteorologic forcing (fields 7-14), information about the rectilinear model grid (field 15), initial conditions (field 16), and the recharge forcing applied at the surface (field 17).
The process-based hydrologic model receives additional input information not included in Table 2.5, such as cell dimensions and time step sizes. However, we omit many of these fields from our input features because they are temporally invariant and constant across all training examples. Thus, each surrogate architecture likely learns this information implicitly during training. Some fields, such as cell volume and the initial pressure head for Stage 1 (§2.2), are constant across all simulations, but are included nonetheless to support future studies seeking to modify these parameters.
@@ 2.5.2 Broadcasting
Broadcasting streamlines manipulation of inputs and targets through the use of large, combined tensors. The input features (Table 2.5) include fields that remain constant in time, fields that are constant in space, and fields that are constant in both space and time. Broadcasting simply repeats the values of each field across dimensions where that field remains constant, which ensures that the tensor for each input feature is four-dimensional (x, y, z, and time). For example, we broadcast air temperature at each time step — which does not vary spatially across the domain — along the x, y, and z dimensions. Similarly, given that saturated hydraulic conductivity does not change from one time step to the next, we repeat the 3D hydraulic conductivity field from the first time step across all time steps. This approach standardizes the dimensions of all input fields and allows them to be combined into one large tensor for input into each architecture.
@@ 2.5.3 Downsampling
While broadcasting standardizes input dimensions, it increases the total size of the input features and, by extension, the memory allocated to each training example. The memory required to train an ML model is an important component of computational efficiency; if the input and output for a single training example exceeds the memory capacity of a single graphics processing unit (GPU), then the model will have to be trained across multiple GPUs, which will increase the total number of processor-hours used during training. Modern hydrologic model domains can include anywhere from $10^{5}$ to $>$$10^{8}$ total grid cells (Wood et al., 2011; Bierkens et al., 2015). At this scale, after broadcasting each input feature across the x, y, z, and time dimensions and adding in the size of the training targets (i.e., the transient 3D pressure fields from ParFlow-CLM), the memory allocated to a single training example can quickly exceed 100 GB, which is larger than the memory capacity of most modern GPUs. To address this challenge, we limit the size of a single training example through temporal and spatial downsampling.
For Stages 1 and 2 (§2.2), we apply temporal downsampling. These model stages are only used to initialize water content throughout the vadose zone. Because intermediate pressure fields are not used to calculate recharge efficiency, decreasing the temporal resolution does not limit downstream analyses. For each stage, we sample 16 output time steps from the spin up period using a geometric sampling scheme:
$\left\{\left\lfloor\sum_{k=1}^{t}r^{k-1}\right\rfloor:\sum_{k=1}^{16}r^{k-1}=160\right\}_{t=1}^{16}$ (1)
The length of 16 is chosen so that a single training example fits on our training hardware, while the geometric sampling scheme prioritizes time step outputs toward the beginning of the model stage, when changes in pressure head are most pronounced. This flexibility in time step outputs emphasizes the advantage of generating simulation outputs with ML models: while a process-based model may require short time steps to converge to a solution, an ML model can reasonably learn patterns across larger time scales and can take arbitrarily large time steps. Though the final steady-state pressure field at the end of each stage is the only output used in subsequent modeling stages, we calculate loss across multiple time steps from this stage to ensure that the model simulates physically meaningful processes (i.e., gradual changes in pressure over time). This also forces the model to output intermediate pressure fields; visually examining these intermediate outputs allows the user to more easily diagnose factors contributing to poor model performance. While this geometric sampling scheme results in a variable time step size within an individual simulation, the interval between sampled time steps is consistent across all training examples. Thus, the data-driven ML models should implicitly learn these temporal patterns during training.
@@
To downsample the Stage 3 training targets, we use an autoencoder. The autoencoder is a separate neural network that is trained to reproduce the spatially variable pressure field from a single time step, but it is forced to store this information in an intermediate, lower-dimensional space (Figure 3(a)). Each surrogate model is then trained on the lower-dimensional representation of the 3D pressure fields, known as the encoded targets (Figure 3(b)). We train a single autoencoder that is used for all surrogate models. To calculate surrogate performance on the test set, we then decode the surrogate model’s predictions using the autoencoder and calculate error metrics relative to the high-resolution pressure fields output by ParFlow-CLM. This workflow allows us to train our surrogate model in the reduced dimensional space, where tensors are 4% their original size, while retaining the high spatial resolution of our predictions for use in downstream tasks, such as calculating recharge efficiency. For a detailed overview of the autoencoder architecture, please refer to Section S1.3 in the Supplementary material.
@@ 2.6.1 Training configuration
where $y_{i}$ is the ground-truth output for the $i$th training example, $x_{i}$ is the input for the $i$th training example, $h(x_{i})$ is the predicted output, $w$ defines a depth-wise weighting scheme, and $||\cdot||_{2}$ is the L2 norm. We choose this normalized loss function because it is particularly effective when there is large variance in the data and because previous hydrologic surrogate modeling work suggests that it provides faster convergence than the L1 loss (Wen et al., 2022). Given that managed aquifer recharge efficiency is largely governed by unsaturated flow processes (Perzan et al., 2023a), accurately simulating pressure changes within the vadose zone is more important than simulating pressure changes deep within the saturated zone. Thus, we define the weights $w$ such that they more severely penalize pressure errors closer to the surface than at depth (Figure S3).
To investigate the impact of our choice of loss function on model performance, we train additional CNN3d surrogates on Stage 1 (§2.2) simulations using the mean square error loss (i.e., the square of the numerator of Equation 2). Due to computational constraints, we could not train duplicates of each surrogate architecture using the mean square error loss, so we only perform this comparison with the CNN3d because it has the fastest training time.
We perform a random 80%-10%-10% (217-28-28) training–validation–test split and all performance metrics are reported based on the test set unless otherwise noted. Features and labels are normalized using their respective minimum and maximum values from the training set. We normalize all splits by the training set’s minimums and maximums to ensure the model does not inadvertently gain knowledge of the validation and test sets before it is supposed to evaluate on them. To assess the effects of data normalization on ML model accuracy, we train two additional Stage 1 CNN3d models: one using Z-score normalization and one without any data normalization. As discussed above, we only perform these tests using the CNN3d because of its short training time.
@@ 2.6.2 Hardware environment
We train all ML models on a single NVIDIA A100 Tensor Core graphics processing unit (GPU) with 80 GB memory. All ParFlow-CLM simulations were performed on AMD Epyc 7543 central processing units (CPUs). Depending on the model stage (§2.2), ParFlow-CLM simulations were run either on 4 MPI ranks with 1 CPU core per rank (Stage 1 and Stage 3) or 6 MPI ranks with 1 CPU core per rank (Stage 2).
@@ 2.6.3 Model accuracy evaluation
We evaluate each surrogate model using the mean absolute percentage error (MAPE) between the predicted pressure head time-series and the ground-truth (ParFlow-CLM) pressure head time-series in a cell-by-cell manner. Because we train a different surrogate for each modeling stage, we report different MAPE values for each stage. As with the loss function, we weight MAPE by depth to ensure that each model’s performance accurately reflects its ability to simulate vadose zone flow (Figure S3). We also evaluate model performance when each surrogate is run sequentially across all three stages (§2.2), in what we call an end-to-end (E2E) evaluation. The only difference between a simulation generated during training and a simulation generated in an E2E manner is that, during training, the initial pressure head fields (field 16 in Table 2.5) of Stages 2 and 3 are populated with the ground-truth initial pressure head field produced by ParFlow-CLM. During E2E evaluation, the initial pressure head fields of Stages 2 and 3 are populated with the final pressure head output of Stages 1 and 2, respectively. E2E provides a more realistic evaluation of surrogate performance in a real-world application.
Training an accurate ML surrogate requires a sufficiently large training set, but at a certain point, the cost of increasing the training set size may outweigh any benefit in improved prediction performance. To evaluate this relationship and estimate the minimum number of process-based simulations required to train an accurate surrogate, we train multiple versions of each Stage 1 surrogate model while using smaller and smaller training sets. We only consider models trained on Stage 1 simulations in this section to simplify analysis and minimize additional training. In addition to the baseline training set size of 245 examples used for training (with a 217-28 training-validation split), we use training set sizes of 196 (174-22 split), 147 (130-17), 98 (87-11), 49 (43-6), 25 (22-3), 13 (11-2), 8 (7-1), 5 (4-1), and 3 (2-1) training examples while maintaining the test set constant at 28 examples. Because ML models typically exhibit a power-law relationship between model error and training set size (Hestness et al., 2017), we then fit a power function to these data ($y=am^{b}$, where $m$ is the number of training examples, $y$ is the test set MAPE, and $a$ and $b$ are regression coefficients). Lastly, we differentiate $am^{b}$ with respect to $m$, which yields $abm^{b-1}$. This expression represents the instantaneous change in MAPE given an increase in training set size, which we use to compare each model’s sensitivity to training set size.
@@ 2.6.4 Model runtime evaluation
In addition to evaluating ML model accuracy, we quantify the total compute time needed to generate simulations with this hybrid modeling workflow. Since model runtime is largely dependent on the hardware on which it is executed, we strictly adhere to the hardware environment described in §2.6.2 when evaluating model runtime. To simplify analysis and minimize additional training, we only consider models trained on Stage 1 (§2.2) simulations in this section.
@@ 2.7 Summary of experiments
    We train all seven surrogate architectures — the CNN4d, ViT4d, U-FNO4d, CNN3d, ViT3d, U-FNO3d, and PredRNN++ (§2.4) — on Stage 1, 2, and 3 ParFlow-CLM simulations (§2.2), amounting to 21 trained models. We also train an autoencoder for all Stage 3 models. This totals 22 trained models.
@@ 3.1 Model accuracy
All seven ML models achieve $<$10% MAPE on the test set (Figure 5 and Table S8), which is the threshold for “highly accurate forecasting” (Lewis, 1982). The PredRNN++ displays the lowest error for Stage 1 and end-to-end (E2E) predictions, while the CNN3d and ViT3d perform the best on Stage 2 and Stage 3, respectively. Even though the non-sequential models have more parameters than their one-step counterparts (Table 1), they do not exhibit an appreciable increase in predictive performance. This result suggests that model size alone is not a reliable indicator for model performance and the type of architecture (one-step vs. non-sequential) may have a stronger impact on model performance. One architectural property that lends an advantage for spatiotemporal predictions is the use of large memory cells in a recurrent manner by the PredRNN++. Indeed, the PredRNN++ outperforms the others in E2E evaluation, the measurement most representative of real-world applications. The large memory cells retained between time steps likely provide ample degrees of freedom for the model to store spatially heterogeneous information over time, and these information stores are referenced recurrently, honestly reflecting the step-by-step manner in which the simulations are generated by the ground-truth ParFlow-CLM.
Figure 5: Test set MAPEs for all ML models across all stages. (b) is a local enlargement of (a), indicated by the “Inset” label in (a). The ML surrogate models are grouped by model type (see §2.4), with the non-sequential models on the left (CNN4d, ViT4d and U-FNO4d), the one-step models in the middle (CNN3d, ViT3d and U-FNO3d) and the recurrent model on the right (PredRNN++). Exact MAPE values are listed in Table S8.
Overall, prediction error is the highest for Stage 1, with a mean MAPE of 5.6% across all ML models, as opposed to 0.13% and 0.42% for Stages 2 and 3, respectively. The higher Stage 1 MAPE could result from its large time scale (§2.2 and 2.5.3). While Stages 2 and 3 span 1 and 2 years, respectively, Stage 1 spans over 400 years, with more dramatic changes in pressure between outputs. Pressure head changes across large time steps may be more difficult to learn than smaller, more incremental changes, potentially contributing to the poorer performance of Stage 1 models. Sampling the spin-up period at higher temporal resolution (§2.5.3) may improve surrogate model accuracy.
The higher MAPE in Stage 1 could also be explained by the fact that its pressure head field is initialized not with a ground-truth pressure head field but with pressure values in hydrostatic equilibrium with the water table at 45 m depth (Table 2.5). By contrast, the pressure heads of Stages 2 and 3 are initialized with the ground-truth pressure head field output by ParFlow-CLM at the end of Stages 1 and 2, respectively. Previous modeling work at this site has shown that managed aquifer recharge efficiency is most strongly controlled by vadose zone water content prior to inundation (Perzan et al., 2023a). Thus, stages that use ground-truth initial pressure head fields as inputs (Stages 2 and 3) achieve superior performance than the stage that does not (Stage 1).